Calculate the mass concentration of the solution if there is 0.12 mol KCl in 500 cm3 of solution!


1
Expert's answer
2021-05-13T02:08:54-0400

CM = n / V

n = CM x V = 0.12 x 0.5 = 0.06 mol

n = m / M

M (KCl) = 74.6 g/mol

m (KCl) = n x M = 0.06 x 74.6 = 4.48 g

Lets assume that KCl solution density is 1 g/ml. Therefore, m (KCl)solution = 500 x 1 = 500 g.

% (KCl) = (100 x 4.48) / 500 = 0.9%